A body has an initial velocity of 88m/sec moving with an acceleration retardation of 2 m/ sec2 find the time after which it will stop and the distance it will travel before coming at rest

Grade:12th pass

1 Answers

Saiyam Leekha
6 years ago
Given: u = 88m/s
a = -2m/s, v = 0m/s
V2-u2=2as
0- 88= (2)(-2)s
: s = 1936m
And v = u + at    
     0 = 88 + (-2)t
     t = 44s